# TFB: Towards Comprehensive and Fair Benchmarking of Time Series Forecasting Methods

### v3

@@ 2. RELATED WORK
Early proposals primarily employed statistical learning methods. ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), Theta (Federico Garza, 2022), VAR (Godahewa et al., 2021), and Kalman Filter (KF) (Harvey, 1990) are classical and widely utilized methods. These forecasting methods are based on the idea that future values of time series can be predicted from observed, past values. With the rapid advances in machine learning technologies, machine learning methods for TSF have emerged (Fischer et al., 2020). Notably, XGBoost (Chen and Guestrin, 2016; Zhang et al., 2021), Gradient Boosting Regression Trees (GBRT) (Friedman, 2001), Random Forests(Breiman, 2001; Mei et al., 2014) and LightGBM (Ke et al., 2017) have been applied extensively to better accommodate nonlinear relationships and complex patterns. Machine learning methods are flexibile at handling different types and lengths of time series and generally offer better forecasting accuracy than traditional methods.
@@ 3. Preliminaries
[…]
@@
2:$Z_{\min}\leftarrow\min(Z),~{}Z_{\max}\leftarrow\max(Z)$
3:$S=\{s_{i}~{}|~{}s_{i}\leftarrow Z_{\min}+(i-1)\frac{Z_{\max}-Z_{\min}}{m},1\leq i\leq m\}$ where m is the number of thresholds
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 4.1. Datasets
METR-LA (Li et al., 2017) Traffic 5 mins 34,272 207 7:1:2
PEMS-BAY (Li et al., 2017) Traffic 5 mins 52,116 325 7:1:2
PEMS04 (Song et al., 2020) Traffic 5 mins 16,992 307 6:2:2
PEMS08 (Song et al., 2020) Traffic 5 mins 17,856 170 6:2:2
Traffic (Wu et al., 2021a) Traffic 1 hour 17,544 862 7:1:2
ETTh1 (Zhou et al., 2021) Electricity 1 hour 14,400 7 6:2:2
ETTh2 (Zhou et al., 2021) Electricity 1 hour 14,400 7 6:2:2
ETTm1 (Zhou et al., 2021) Electricity 15 mins 57,600 7 6:2:2
ETTm2 (Zhou et al., 2021) Electricity 15 mins 57,600 7 6:2:2
Electricity (Trindade, 2015) Electricity 1 hour 26,304 321 7:1:2
Solar (Lai et al., 2018) Energy 10 mins 52,560 137 6:2:2
Wind (Li et al., 2022) Energy 15 mins 48,673 7 7:1:2
Weather (Wu et al., 2021a) Environment 10 mins 52,696 21 7:1:2
AQShunyi (Zhang et al., 2017) Environment 1 hour 35,064 11 6:2:2
AQWan (Zhang et al., 2017) Environment 1 hour 35,064 11 6:2:2
ZafNoo (Poyatos et al., 2020) Nature 30 mins 19,225 11 7:1:2
CzeLan (Poyatos et al., 2020) Nature 30 mins 19,934 11 7:1:2
FRED-MD (McCracken and Ng, 2016) Economic 1 month 728 107 7:1:2
Exchange (Lai et al., 2018) Economic 1 day 7,588 8 7:1:2
NASDAQ (Feng et al., 2019) Stock 1 day 1,244 5 7:1:2
NYSE (Feng et al., 2019) Stock 1 day 1,243 5 7:1:2
NN5 (Taieb et al., 2012) Banking 1 day 791 111 7:1:2
ILI (Wu et al., 2021a) Health 1 week 966 7 7:1:2
Covid-19 (Panagopoulos et al., 2021) Health 1 day 1,392 948 7:1:2
Wike2000 (Gasthaus et al., 2019) Web 1 day 792 2,000 7:1:2
@@ 4.3.1. Evaluation strategies
Fixed forecasting. Given a time series on length $n$, $f$ future time points are predicted from $n-f$ historical time points, which is illustrated in Figure 6(a).
@@ 4.3.2. Evaluation metrics
[…]

### v4

@@ 1. Introduction
  • We have also launched an online time series leaderboard: https://decisionintelligence.github.io/OpenTS/OpenTS-Bench/.
@@ 2. RELATED WORK
Early proposals primarily employed statistical learning methods. ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), Theta (Federico Garza, 2022), VAR (Toda and Phillips, 1994), and Kalman Filter (KF) (Harvey, 1990) are classical and widely utilized methods. These forecasting methods are based on the idea that future values of time series can be predicted from observed, past values. With the rapid advances in machine learning technologies, machine learning methods for TSF have emerged (Fischer et al., 2020). Notably, XGBoost (Chen and Guestrin, 2016; Zhang et al., 2021), Gradient Boosting Regression Trees (GBRT) (Friedman, 2001), Random Forests(Breiman, 2001; Mei et al., 2014) and LightGBM (Ke et al., 2017) have been applied extensively to better accommodate nonlinear relationships and complex patterns. Machine learning methods are flexibile at handling different types and lengths of time series and generally offer better forecasting accuracy than traditional methods.
@@ 3. Preliminaries
(1) $\footnotesize\mathit{Trend\_Strength}=\mathit{max}\left(0,1-\frac{\mathit{var}\left(R\right)}{\mathit{var}\left(X-S\right)}\right),$
@@
[…]
@@
8:return $\text{$\delta$}\leftarrow\mathit{median}(\{M_{1}^{\prime},M_{2}^{\prime},...,M_{m}^{\prime}\})$
@@
4:Define index $I=\mathit{argsort}(Y)\in\mathbb{R}^{T^{\prime}\times 1}$, then characterize $Y$ to obtain $Z\in\mathbb{R}^{T^{\prime}\times 1}$:
@@
(2) $\footnotesize\mathit{Seasonality\_Strength}=\max\left(0,1-\frac{\mathit{var}\left(R\right)}{\mathit{var}\left(X-T\right)}\right)$
@@
Stationarity refers to the mean of any observation in a time series $X=\langle x_{1},x_{2},...,x_{n}\rangle$ is constant, and the variance is finite for all observations. Also, the covariance $\mathit{cov}(x_{i},x_{j})$ between any two observations $x_{i}$ and $x_{j}$ depends only on their distance |ji||j-i|, i.e., $\forall i+r\leq n,j+r\leq n$ $(\mathit{cov}(x_{i},x_{j})=\mathit{cov}(x_{i+r},x_{j+r})).$ Strictly stationary time series are rare in practice. Therefore, weak stationarity conditions are commonly applied (Lee, 2017)  (Nason, 2006). In our paper, we also exclusively focus on weak stationarity.
@@
(3) $\footnotesize\mathit{Stationarity}=\begin{cases}\mathit{True},&\mathit{ADF}(X)\leq 0.05\\ \mathit{False},&\mathit{ADF}(X)>0.05\end{cases}$
@@
(6) $\footnotesize\mathit{Correlation}=\mathit{mean}\left(P\right)+\frac{1}{1+\mathit{var}\left(P\right)}$
@@ 4.1. Datasets
METR-LA Traffic 5 mins 34,272 207 7:1:2
PEMS-BAY Traffic 5 mins 52,116 325 7:1:2
PEMS04 Traffic 5 mins 16,992 307 6:2:2
PEMS08 Traffic 5 mins 17,856 170 6:2:2
Traffic Traffic 1 hour 17,544 862 7:1:2
ETTh1 Electricity 1 hour 14,400 7 6:2:2
ETTh2 Electricity 1 hour 14,400 7 6:2:2
ETTm1 Electricity 15 mins 57,600 7 6:2:2
ETTm2 Electricity 15 mins 57,600 7 6:2:2
Electricity Electricity 1 hour 26,304 321 7:1:2
Solar Energy 10 mins 52,560 137 6:2:2
Wind Energy 15 mins 48,673 7 7:1:2
Weather Environment 10 mins 52,696 21 7:1:2
AQShunyi Environment 1 hour 35,064 11 6:2:2
AQWan Environment 1 hour 35,064 11 6:2:2
ZafNoo Nature 30 mins 19,225 11 7:1:2
CzeLan Nature 30 mins 19,934 11 7:1:2
FRED-MD Economic 1 month 728 107 7:1:2
Exchange Economic 1 day 7,588 8 7:1:2
NASDAQ Stock 1 day 1,244 5 7:1:2
NYSE Stock 1 day 1,243 5 7:1:2
NN5 Banking 1 day 791 111 7:1:2
ILI Health 1 week 966 7 7:1:2
Covid-19 Health 1 day 1,392 948 7:1:2
Wike2000 Web 1 day 792 2,000 7:1:2
@@ 4.3.1. Evaluation strategies
Fixed forecasting. Given a time series on length $n$, $f$ future time points are predicted from nfn-f historical time points, which is illustrated in Figure 6(a).
@@ 4.3.2. Evaluation metrics
(7) $\mathit{MAE}=\frac{1}{h}\sum_{k=1}^{h}|F_{k}-Y_{k}|$
(9) $\mathit{MSE}=\frac{1}{h}\sum_{k=1}^{h}(F_{k}-Y_{k})^{2}$
(11) $\mathit{RMSE}=\sqrt{\frac{1}{h}\sum_{k=1}^{h}(F_{k}-Y_{k})^{2}}$
(8) $\mathit{MAPE}=\frac{1}{h}\sum_{k=1}^{h}\frac{|Y_{k}-F_{k}|}{Y_{k}}\times 100\%$
(10) $\mathit{SMAPE}=\frac{100\%}{h}\sum_{k=1}^{h}\frac{|F_{k}-Y_{k}|}{\frac{|Y_{k}|+|F_{k}|}{2}}$
(12) $\mathit{WAPE}=\frac{\sum_{k=1}^{h}|Y_{k}-F_{k}|}{\sum_{k=1}^{h}|Y_{k}|}$
(13) $\footnotesize\mathit{MSMAPE}=\frac{100\%}{h}\sum_{k=1}^{h}\frac{|F_{k}-Y_{k}|}{\max(|Y_{k}|+|F_{k}|+\epsilon,0.5+\epsilon)/2}$
(14) $\footnotesize\mathit{MASE}=\frac{\sum_{k=M+1}^{M+h}|F_{k}-Y_{k}|}{\frac{h}{M-S}\sum_{k=S+1}^{M}|Y_{k}-Y_{k-S}|},$